OpenStudy (anonymous):

sqrt -48 simplify the following expression express in terms of i

OpenStudy (anonymous):

sqrt(xy) = sqrt(x)*sqrt(y) i = sqrt(-1) Any guesses what to do next?

OpenStudy (anonymous):

Multiply

OpenStudy (anonymous):

Multiply what? use the info from the problem.

OpenStudy (anonymous):

Nm it would be to divide right?

OpenStudy (anonymous):

bc when I tried finding the sqrt it was a decimal

OpenStudy (anonymous):

You can't use a calculator to solve this problem. What you need to do is factor out the square root. sqrt(28) = 2*sqrt(7) because sqrt (28) is equal to sqrt(4)*sqrt(7) = sqrt(4*7) = sqrt(28)

OpenStudy (anonymous):

Or you would do (-1)6*8=48

OpenStudy (anonymous):

Yes, you're getting closer :)

OpenStudy (anonymous):

Ugh I am stuck

OpenStudy (anonymous):

So would I then break down the 8?

OpenStudy (anonymous):

and the 6?

OpenStudy (anonymous):

You're not stuck, you are so close. You have to factor out numbers that can be broken down by square roots. Like in my example of sqrt(28), I pulled out the 4 because sqrt(4) = 2

OpenStudy (anonymous):

You can't get a sqrt of 6 or 8. But what factors of 48 CAN be broken down by a square root?

OpenStudy (anonymous):

4?

OpenStudy (anonymous):

Yes, that's one. But there is a greater factor...

OpenStudy (anonymous):

12

OpenStudy (anonymous):

Yes, but you can't get a sqrt of 12... a little higher Think only in terms of numbers that are squares...

OpenStudy (anonymous):

16

OpenStudy (anonymous):

YES! :)

OpenStudy (anonymous):

Yay! So how can I write it out??

OpenStudy (anonymous):

So, 16 is the biggest factor of 48 that will help us here. 16*3 = 48 -1*16*3 = -48 Here's our problem, what is sqrt(-48) Well, sqrt(-48) = sqrt(-1*16*3) What's next?

OpenStudy (anonymous):

-1*8*3^2?

OpenStudy (anonymous):

Not quite. Use this rule, sqrt(a*b) = sqrt(a)*sqrt(b)

OpenStudy (anonymous):

16*3?

OpenStudy (anonymous):

Ohh!!!! Ok hang on!

OpenStudy (anonymous):

Ok so I got a big #

OpenStudy (anonymous):

2304 which doesnt look right

OpenStudy (anonymous):

No that's not correct. If sqrt(a*b) = sqrt(a)*sqrt(b), what does sqrt(-1*16*3) = ???

OpenStudy (anonymous):

-48

OpenStudy (anonymous):

Sorry I told you im not good at this ;/

OpenStudy (anonymous):

It's ok. You are trying to pull the -1, 16, and 3 out of the sqrt. If sqrt(a*b) = sqrt(a)*sqrt(b), then for example sqrt(4*7) = sqrt(4)*sqrt(7)

OpenStudy (anonymous):

one sqrt becomes two sqrts

OpenStudy (anonymous):

I got 4 for sqrt 16

OpenStudy (anonymous):

Right. But what is the whole thing? If sqrt(a*b*c) = sqrt(a)*sqrt(b)*sqrt(c) then, sqrt(-1*16*3) = sqrt(-1)*sqrt(16)*sqrt(3) right?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

So then that would be 0

OpenStudy (anonymous):

zero? where are you getting zero?

OpenStudy (anonymous):

By putting sqrt after the numbers the calculator zeroed it out

OpenStudy (anonymous):

Put the calculator down :)

OpenStudy (anonymous):

You do not need it at all. Trust me.

OpenStudy (anonymous):

LOL ok

OpenStudy (anonymous):

Ok, so we know that sqrt(-48) = sqrt(-1)*sqrt(16)*sqrt(3) right?

OpenStudy (anonymous):

Right. So do we add or subtract at all?

OpenStudy (anonymous):

There is no addition or subtraction in this problem. Don't look for things that aren't there.

OpenStudy (anonymous):

What you have is sqrt(-48) = sqrt(-1)*sqrt(16)*sqrt(3) Simplify.

OpenStudy (anonymous):

[You already told me sqrt(16) is 4, so 4*sqrt(-1)*sqrt(3)]

OpenStudy (anonymous):

so simplifying 4 would be 2

OpenStudy (anonymous):

4 is already simplified. 4 is the sqrt of 16

OpenStudy (anonymous):

you can't make 4 simpler than 4 :)

OpenStudy (anonymous):

Oh ok oops

OpenStudy (anonymous):

So it would then be sqrt 4 right?

OpenStudy (anonymous):

It's ok, this isn't simple stuff. There's only two things left that might be able to be simplified: sqrt(-1) and sqrt(3) Can you simplify one of these? both?

OpenStudy (anonymous):

No 4 is already simplified. Ignore 4. that's done.

OpenStudy (anonymous):

You can't simplify 3 or -1

OpenStudy (anonymous):

You are half right. You cannot simplify sqrt(3) But, what is sqrt(-1) ???

OpenStudy (anonymous):

1! ;)

OpenStudy (anonymous):

No, that's not correct. A square root is a number when multiplied by itself equals the square. So basically, you are saying that 1 times 1 equals -1.

OpenStudy (anonymous):

Let me ask you this question... what is i ??

OpenStudy (anonymous):

sqrt of -1 =i

OpenStudy (anonymous):

Exactly. So what is the sqrt(-1) ???

OpenStudy (anonymous):

i

OpenStudy (anonymous):

Perfect :)

OpenStudy (anonymous):

So the answer is 4i

OpenStudy (anonymous):

What happened to that sqrt(3) ???

OpenStudy (anonymous):

Don't forget it.

OpenStudy (anonymous):

So, hears your answer: sqrt(-48) = sqrt(16*-1*3) = sqrt(16)*sqrt(-1)*sqrt(3) = 4i*sqrt(3)

OpenStudy (anonymous):

4i*sqrt(3) is the final answer

OpenStudy (anonymous):

I got it!!! Thank you soooo much!!!

OpenStudy (anonymous):

You're welcome :)